OpenStudy (anonymous):

PLEASE HELP! Use elimination to solve the system of equations -3x+ y + 7z = -20 5x - 2y - z = 34 -x + y +4z = -8

OpenStudy (jdoe0001):

have you done elimination on a system of equations with 2 variables?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so the way you'd 3 variables is by picking 2 first, eliminate one pick another 2, eliminate the same one then solve the 2 resulting two-variable functions -3x+ y + 7z = -20 <--- let's use this one 5x - 2y - z = 34 <--- and this one -x + y +4z = -8 \(\bf \begin{array}{llll} -3x+ y + 7z = -20&&-3x+ y + 7z = -20\\ \quad 5x - 2y - z = 34&\times 7\implies &35x-14y-7z=238\\ \hline\\ &&32x-13y+0z=218\\ \quad \\ \color{blue}{32x-13y=218} \end{array}\)

OpenStudy (jdoe0001):

so for our next pair of functions, let us use -3x+ y + 7z = -20 5x - 2y - z = 34 <--- this one -x + y +4z = -8 <--- and this one so \( \begin{array}{llll} \quad 5x - 2y - z = 34&\times 4\implies &20x-8y-4z=136\\ -x + y +4z = -8&&-x + y +4z = -8\\ \hline\\ &&19x-7y+0z=128\\ \quad \\ \color{blue}{19x-7y=128} \end{array}\)

OpenStudy (jdoe0001):

so once you have those 2 RESULTANT functions, you can solve the system of 2 variables by elimination or substitution to get both variables once you have "x" and "y", you can plug those in on any of the 3 original functions to get "z"

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw